A) A sample of hydrogen gas in a rigid container has a pressure of 4.00 atmat 127 °C. At what temperature will the pressure decrease to 190 torr?

B) 0.200 g magnesium metal is completely reacted with excess hydrochloric acid to produce hydrogen gas and magnesium chloride. What volume (in mL) of hydrogen gas could be collected from the reaction at 1.0 atm and 25 °C?
1
Expert's answer
2020-11-26T05:22:33-0500

A) initial, pressure, P1 = 4.00 atm

= 4×760 torr

= 3040 torr

Temperature = 127 °C

So, T1 = 400 K

Final, pressure, P2 = 190 torr

Temperature,T2 = ?

As we know, P1/T1 = P2/T2

Or, T2 = (P2/P1)×T1

Or, T2 = (190/3040)× 400

= 25 K

So,

At 25 K temperaturethe pressure decrease to 190 torr



B) Molar mass of magnesium = 24.3 g/mol

So, 0.200 g magnesium = 8.23×10^-3 mole


Reaction of magnesium with Hydrochloric acid is,

Mg + 2HCl -------> MgCl2 + H2

1 mole of magnesium produces 1 mole H2

8.23×10^-3 mole magnesium produces

= 8.23×10^-3 mole of H2 gas


So,

number of moles of gas, n = 8.23×10^-3 mole

Temperature, = 25°C

T = 298 K

Pressure, P = 1 atm


We know, PV = nRT

Or, V = nRT/P

Or, V = 8.23×10^-3×0.082×298/1

= 0.2011 L

= 201.1 ml

So,

201.1 ml of hydrogen gas could be collected from the reaction at 1.0 atm and 25 °C